1. A typical student listening attentively to a physics lecture class has a heat output of 100 W. How much heat energy does a class of 90 physics students release into a lecture hall over the course of a 50-min lecture?
  2. Assume that all the heat energy in part 1 is transferred to the 3200 m3 of air in the room. The air has a specific heat capacity of 1020 J/kg-K and a density of 1.20 kg/m3. If none of the heat escapes and the air conditioning system is off, how much will the temperature of the air in the room rise during the 50-min lecture?
  3. If the class is taking an exam, the heat output per student rises to 280 W. What is the temperature rise during 50 min in this case?
1
Expert's answer
2021-02-05T17:06:27-0500

1.

"90\times100 \frac {J}{s}\times3\times 10^3s=2.7\times 10^7 J"

2.


"\varDelta t=\frac {Q}{cm}"


"\varDelta t=\frac {2.7\times 10^7 J}{1020\frac {J}{kg-K} \times 1.2 \frac {kg}{m^3}\times 3200 m^3}=6.9 K"

3.

"90\times280 \frac {J}{s}\times3\times 10^3s=7.56\times 10^7 J"


"\varDelta t=\frac {7.56\times 10^7 J}{1020\frac {J}{kg-K} \times 1.2 \frac {kg}{m^3}\times 3200 m^3}=19.3 K"